[Brahms-l] Shift report 20040216 08:00-16:00

From: Apache <apache@bnl.gov>
Date: Mon Feb 16 2004 - 07:57:49 EST
------------------------------------------------------------
Date: Mon Feb 16 07:57:49 2004, Shift: 08:00-16:00
Supervisor: BHS, with: 
Log book#15 pages 58 - 58

Shift Summary:
------------------------------------------------------------
1) It's been a rough night for the machine. It had a blue QLI when I came in, but got beam on at 01:00. Then it took over an hour to steer properly (before that our collisions were useless), and the store took about 2.5hrs after this again to fall to below useful rates again. We requested dump with PHENIX, but just as we were preparing for it the beam died anyway due to a cryo problem. This problem is still under investigation, no estimate for beam.

2) Three data runs at 45deg 1050A, 4deg 1/1A:
10384-10386
Tr2: 217k
Tr3: 477k
This brings us almost up to the next MRS milestone of 2M tr3 (now have 1.96M since we started this particular round of 1050A). The FS has approx 1M to go at this setting.

3) No detector problems. One TOFW tube (TOP61) shows a voltage discrepancy of +20V, while all the others are less than 4V from the setpoint - I don't think it did this before.

Also, no sign of mice. Maybe Ionut the Terrible scared them all off last night ;-)
------------------------------------------------------------

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Mon Feb 16 07:57:59 2004

This archive was generated by hypermail 2.1.8 : Mon Feb 16 2004 - 07:58:23 EST